# Evolution of Semi-convective Staircases in Rotating Flows: Consequences for Fuzzy Cores in Giant Planets

### v1

@@ Abstract
Recent observational constraints on the internal structure of Jupiter and Saturn suggest that these planets have “fuzzy" cores, i.e., radial gradients of the concentration of heavy elements that might span $50\%$ to $70\%$ of each planet’s radius. These cores could be composed of a semi-convective staircase, i.e., multiple convective layers separated by diffusive interfaces arising from double-diffusive instabilities. However, to date, no study has demonstrated how such staircases can avoid layer mergers and persist over evolutionary time scales. In fact, previous work has found that these mergers occur rapidly, quickly leading to only a single convective layer. Using 3D simulations of convective staircases in non-rotating and rotating flows, we demonstrate that rotation prolongs the lifetime of a convective staircase by increasing the timescale for both layer merger and erosion of the interface between the final two layers. We present an analytic model for the erosion phase, predicting that rotation increases the erosion time by a factor of approximately $\mathrm{Ro}^{-1/2}$, where $\mathrm{Ro}$ is the Rossby number of the convective flows (the ratio of the rotation period to the convective turnover time). For Jovian conditions at early times after formation (when convection is vigorous enough to mix a large fraction of the planet), we find the erosion time to be roughly $10^{9}~{}\mathrm{yrs}$ in the non-rotating case and $10^{11}~{}\mathrm{yrs}$ in the rotating case. Thus, the current existence of convective staircases within the deep interiors of giant planets is a strong possibility, and rotation could be an important factor in the preservation of their fuzzy cores.
@@ 1 Introduction
Determining whether gas giants are fully mixed is crucial for understanding their internal structure and evolution. The strength and distribution of compositional gradients affect heat transport, long-term cooling, and magnetism. Prior to the Juno and Cassini missions, gas giants were thought to be composed of a central core of heavy elements underneath a convective hydrogen-helium envelope. However, measurements of Jupiter’s gravity field (Bolton et al., 2017a, b; Wahl et al., 2017; Militzer et al., 2022; Howard et al., 2023) and seismology of Saturn’s rings (Mankovich & Fuller, 2021) stand in conflict with that picture, and instead point toward “fuzzy” cores wherein heavy element abundances smoothly decrease from the core into the outer regions (e.g., Helled & Stevenson, 2024).
@@ 2.1 Qualitative evolution of a convective staircase
For example, in the nonrotating model (Figure 2a), the two innermost and the two outermost layers grow in thickness and engulf the layers sandwiched inbetween. This process completes around $t\approx 0.03~{}\tau_{T}$, leaving four layers, two inner layers and two outer layers. The outer layers then start to devour the inner layers with the engulfment requiring $\sim 0.07$ more thermal diffusion times. The penultimate state, two layers of roughly equal size, is achieved around $t\approx 0.1~{}\tau_{T}$. These two final layers persist over time, since there is a density jump across the common interface that prevents their instantaneous merger and mixing. However, the layers slowly mix across the interface and the density difference between them decreases until the final merger at roughly $t=0.2~{}\tau_{T}$. Hence, the erosion of the final interface takes roughly one-tenth of a thermal diffusion time. Note, this erosion process is a turbulent mixing mechanism (i.e., entrainment) not a diffusive one. The compositional diffusion time across the entire box is an order of magnitude larger than the thermal diffusion time ($\kappa_{C}/\kappa_{T}=0.1$, where $\kappa_{C}$ is the chemical diffusivity). Since, the interfaces have a width that is roughly $0.1~{}L_{z}$, the time required for heavy elements to diffuse across the interface is about one thermal diffusion time, which is an order of magnitude longer than the erosion time exhibited by the simulations. The rotating model (Figure 2b), appears to undergo similar processes, except that the mergers and interface erosion take much longer to occur. The initial layer mergers do not complete until $t=0.33~{}\tau_{T}$ and the final internal interface takes roughly 0.2 thermal diffusion times to erode, i.e., two times longer than the non-rotating case.
@@
As we show later, an important parameter for the mixing timescales of the fluid is the Rossby number, defined as the ratio of the rotational period to the convective turnover time, $\mathrm{Ro}\equiv U_{\mathrm{conv}}/2\Omega\ell$, where $U_{\mathrm{conv}}$ is the characteristic scale of the convective velocity, and $\ell$ is the characteristic lengthscale of the convection. Rapidly rotating flows have $\mathrm{Ro}\ll 1$, and the efficiency of convective mixing is much smaller than for non-rotating flows. In our simulations, we calculate Ro using the r.m.s. velocity of the fluid in each convective layer as a proxy for $U_{\mathrm{conv}}$, and set $\ell$ to the thickness of a convective layer in the staircase. Figure 3 shows profiles of the r.m.s. velocity for the rotating case, at different stages during the evolution of the staircase. As expected, we find that as layers merge the convective velocity increases because the freefall distance in each surviving layer increases. Hence, the Rossby number (reported on top of each curve in Figure 3) decreases after each merger, starting with $\mathrm{Ro}\approx 0.107$ in the 6-layers phase, and ultimately reaching $\mathrm{Ro}\approx 0.044$. We also note that initially, when there are 6 layers in the fluid, the velocity and size of each layer is uniform. Later on, the outer convective layers have a larger speed than the interior layers, engulfing them and leaving only two layers that slowly erode each other until a single convective layer remains.
@@ 3 Analytic model for the erosion phase
A complete characterization of all the timescales involved in the problem, and how they change with the choice of parameters is beyond the scope of this work. However, in what follows, we present an analytic model for the erosion stage that reproduce relatively well the numerical results.
Figure 4: Panel (a): Density profile at the beginning of the erosion phase for the non-rotating and rotating cases. Note that in the rotating case, the density is not constant within the convection zone. Instead, there is a positive density gradient due to the thermal structure of the fluid in the convection zone (in nonrotating convection, the temperature becomes isothermal whereas in rotating convection, the fluid develops a small temperature gradient). Panel (b): Evolution of the density constrast between the two layers during the erosion phase. For both cases $\Delta\rho$ decreases linearly with time, but the rotating case decays at a slower rate. Further, the initial density contrast at the beginning of the erosion phase is smaller for the rotating case (likely due to diffusion since it takes longer for the fluid to reach this stage).
@@
where the preceding factor of 2 in the final expression accounts for entrainment from above and below the interface, $\rho_{0}$ is the average density of the fluid, and $\varepsilon$ is the mixing efficiency (0.5absent0.5$\approx$ 0.5≈ 0.51111 for turbulent flows, e.g., Fuentes & Cumming, 2020).
We see that given $U_{\mathrm{conv}}$, Equation (1) can be integrated to obtain $\Delta\rho(t)$. In the non-rotating and rapidly-rotating limits, typical convective speeds ($U_{\rm NR}$ and $U_{\rm R}$, respectively) can be estimated using mixing-length theory and CIA balance—a three way balance between the Coriolis, inertial, and buoyancy force (e.g., Stevenson, 1979; Ingersoll & Pollard, 1982; Barker et al., 2014; Aurnou et al., 2020). Following the derivation of Fuentes et al. (2023), we obtain
@@
where the numbers were inferred from Cumming et al. (2018) and Helled et al. (2022). This is a difference of two orders of magnitude, and the mixing time is longer than the current age of Jupiter or Saturn, thus explaining the presence of fuzzy cores in their interiors. We emphasize clear that evaluating the numbers at later times would give even longer timescales, since convection becomes weaker as the planet’s luminosity decreases with time.
@@ 4 Discussion
To the best of our knowledge, the only study that investigated the influence of rotation on convective staircases under astrophysical conditions was carried by Moll & Garaud (2017). As in this work, they found that rotation always acts to reduce both thermal and compositional transport across layers. However, they did not characterize the mixing timescales or survival of the layers. Interestingly, they found that in some of their simulations where the flow is strongly constrained by rotation (the relevant regime for gas giants) layer formation does not occur, and instead, large-scale vortices form in the fluid, enhancing the transport of material. Unfortunately, their results appear to be highly dependent on the dimensions of the simulation domain likely due to the choice of triply periodic boundary conditions. In this work, we have avoided that by using “fixed-flux” boundary conditions (see the Appendix).
A recent paper by Tulekeyev et al. (2024), where layer formation is investigated in non-rotating flows using fixed-flux boundary conditions, has shown similar results to our simulations, i.e., layer formation followed by several merger events. Initially layers merge by engulfment. This process continues until only two layers remain. The density contrast between these two layers then erodes as entrainment mixes fluid across the separating interface, until a single layer is left.
@@
Despite the simplifications made in our model (e.g., ignoring spherical geometry, density stratification, and magnetic fields) and the uncertainties in the interior parameters of gas giants (particularly for $\Delta\rho_{0}$ and the mixing efficiency $\varepsilon$), we have demonstrated that the effect of rotation is substantial in rapidly rotating gas giants, increasing the erosion time of the staircase by two orders of magnitude, for Jovian conditions at $t\sim 10^{8}~{}\mathrm{yrs}$ after formation. The effect is even larger at later times since the convective strength becomes smaller as the planet’s luminosity decreases with time. Also, a smaller value of the mixing efficiency for the rotating case will just make the timescale longer, once again supporting our hypothesis that rotation prolongs the lifetime of a convective staircase.
@@
We encourage future work on hydrodynamic (and magnetohydrodynamic) models of semiconvection in rapidly rotating flows. This process is not only relevant for the thermal evolution and transport processes in giant planets, but also in stars (e.g., Spruit, 2013; Zaussinger & Spruit, 2013; Ding & Li, 2014; Moore & Garaud, 2016). Recently, semiconvection has been shown to be crucial for understanding the internal dynamics, magnetism, and potential habitability of icy moons (e.g., Vance et al., 2021; Wong et al., 2022; Naseem et al., 2023; Idini & Nimmo, 2024). The Rossby number in icy moons is also small ($\mathrm{Ro}\ll 1$, see, e.g., Bire et al., 2022) and therefore the results obtained in this work could be applied to them. Improvements in numerical modeling by relaxing some of the assumptions above, particularly, simulations in spherical geometry rather than in Cartesian boxes would be of great interest.
@@
We thank Keith Julien for the countless conversations on fluid dynamics, in particular the ones on rotationally-constrained turbulence. You were not only a great scientist, but also a good friend, with an infectious smile. You will be sorely missed. We also thank Pascale Garaud for useful conversations on double-diffusive convection. J.R.F. is supported by NASA through the Solar System Workings grant 80NSSC24K0927, and Heliophysics grants 80NSSC19K0267 and 80NSSC20K0193. B.W.H. is supported by NASA Heliophysics through grants 80NSSC20K0193, 80NSSC24K0125, and 80NSSC24K0271. A.F. acknowledges support from NASA HTMS grant 80NSSC20K1280, and from the George Ellery Hale Postdoctoral Fellowship in Solar, Stellar and Space Physics at The University of Colorado, Boulder. E.H.A. acknowledges support from NSF grant PHY-2309135 and Simons Foundation grant (216179, LB) to the Kavli Institute for Theoretical Physics (KITP). We express the fluid quantities as the sum of a linear hydrostatic background (denoted by the subscript 0) and a dynamic perturbation to the background (denoted by primes), e.g., the total temperature and composition are expressed as $T=T_{0}(z)+T^{\prime}$, and $C=C_{0}(z)+C^{\prime}$, respectively. The density perturbations satisfy $\rho^{\prime}/\rho_{0}\ll 1$, and are related to $T^{\prime}$ and $C^{\prime}$ through $\rho^{\prime}=\rho_{0}(\beta C^{\prime}-\alpha T^{\prime})$, as demanded by the Boussinesq approximation (Spiegel & Veronis, 1960). Here, $\beta$ and $\alpha$ are the coefficients of compositional and thermal contraction/expansion (both assumed positive constants), respectively. Before presenting the fluid equations, we non-dimensionalize them using $[T]=|\partial_{z}T_{0}|L_{z}$, $[C]=|\partial_{z}C_{0}|L_{z}$ as units of temperature and composition. We use the domain’s depth $L_{z}$ as the unit of length, and the thermal diffusion time $\tau_{T}=L^{2}_{z}/\kappa_{T}$ (where $\kappa_{T}$ is the thermal diffusivity) as the unit of time. By this choice, a unit of pressure corresponds to $[P]=\rho_{0}(\kappa_{T}/L_{z})^{2}$. The dimensionless equations are

### v3

@@ Abstract
Recent observational constraints on the internal structure of Jupiter and Saturn suggest that these planets have “fuzzy" cores, i.e., gradients of the concentration of heavy elements that might span a large fraction of the planet’s radius. These cores could be composed of a semi-convective staircase, i.e., multiple convective layers separated by diffusive interfaces arising from double-diffusive instabilities. However, to date, no study has demonstrated how such staircases can avoid layer mergers and persist over evolutionary time scales. In fact, previous work has found that these mergers occur rapidly, leading to only a single convective layer. Using 3D simulations, we demonstrate that rotation prolongs the lifetime of a convective staircase by increasing the timescale for both layer merger and erosion of the interface between the final two layers. We present an analytic model for the erosion phase, predicting that rotation increases the erosion time by a factor of approximately $\mathrm{Ro}^{-1/2}$, where $\mathrm{Ro}$ is the Rossby number of the convective flows (the ratio of the rotation period to the convective turnover time). For Jovian conditions at early times after formation (when convection is vigorous enough to mix a large fraction of the planet), we find the erosion time to be roughly $10^{9}~{}\mathrm{yrs}$ in the non-rotating case and $10^{11}~{}\mathrm{yrs}$ in the rotating case. If these timescales are confirmed with a larger suite of numerical simulations, the existence of convective staircases within the deep interiors of giant planets is a strong possibility, and rotation could be an important factor in the preservation of their fuzzy cores.
@@ 1 Introduction
Determining whether gas giants are fully mixed is crucial for understanding their internal structure and evolution. The strength and distribution of compositional gradients affect heat transport, long-term cooling, and magnetism. Prior to the Juno and Cassini missions, gas giants were thought to be composed of a central core of heavy elements underneath a convective hydrogen-helium envelope. However, measurements of Jupiter’s gravity field (Bolton et al., 2017a, b; Wahl et al., 2017; Militzer et al., 2022; Howard et al., 2023) and seismology of Saturn’s interior with waves in Saturn’s rings (Mankovich & Fuller, 2021) stand in conflict with that picture, and instead point toward “fuzzy” cores wherein heavy element abundances smoothly decrease from the core into the outer regions (e.g., Helled & Stevenson, 2024).
@@ 2.1 Qualitative evolution of a convective staircase
For example, in the nonrotating model (Figure 2a), the two innermost and the two outermost layers grow in thickness and engulf the layers sandwiched in between. This process completes around $t\approx 0.03~{}\tau_{T}$, leaving four layers, two inner layers and two outer layers. The outer layers then start to devour the inner layers with the engulfment requiring $\sim 0.07$ more thermal diffusion times. The penultimate state, two layers of roughly equal size, is achieved around $t\approx 0.1~{}\tau_{T}$. These two final layers persist over time, since there is a density jump across the common interface that prevents their instantaneous merger and mixing. However, the layers slowly mix across the interface and the density difference between them decreases until the final merger at roughly $t=0.2~{}\tau_{T}$. Hence, the erosion of the final interface takes roughly one-tenth of a thermal diffusion time. The rotating model (Figure 2b), appears to undergo similar processes, except that the mergers and interface erosion take much longer to occur. The initial layer mergers do not complete until $t=0.33~{}\tau_{T}$ and the final internal interface takes roughly 0.2 thermal diffusion times to erode, i.e., the merging of the initial layers and the erosion of the final two layers take approximately three times and two times longer, respectively, than in the non-rotating case.
Figure 3: Vertical profiles of the convective and diffusive composition fluxes for the non-rotating case (blue) and rotating case (orange). The fluxes are shown at the beginning of the erosion phase of the remaining two layers. The solid and dashed lines denote the contribution to the flux purely due to convection and diffusion, respectively.
We emphasize that the erosion process is a turbulent mixing mechanism (i.e., entrainment) not a diffusive one. As show in Figure 3, the compositional transport across the interface is mostly dominated by the convective flux in both the rotating and the non-rotating simulations. Certainly, in the outer portions of the interface, turbulent mixing dominates. Even within the very center of the interface, diffusion only carries roughly 50–60% of the composition flux with the remainder carried by turbulent mixing. So while, diffusion does play a role in chemical transport across the interface, it does so primarily by controlling the width of the interface. As shown by Fernando (1989), in a two-layer system, the equilibrium thickness of the diffusive interface is governed by a balance between thickening due to diffusion and thinning caused by convective entrainment. It is also worth mentioning that Under Jovian conditions, the diffusivity ratio is smaller by an order of magnitude with respect to our simulations (i.e. $\kappa_{C}/\kappa_{T}\sim 0.01$). Adopting a smaller chemical diffusivity would further diminish the role of compositional diffusion leading to its subdominance in interface erosion.
@@
As we show later, an important parameter for the mixing timescales of the fluid is the Rossby number, defined as the ratio of the rotational period to the convective turnover time, $\mathrm{Ro}\equiv U_{\mathrm{conv}}/2\Omega\ell$, where $U_{\mathrm{conv}}$ is the characteristic scale of the convective velocity, and $\ell$ is the characteristic lengthscale of the convection. Rapidly rotating flows have $\mathrm{Ro}\ll 1$, and the efficiency of convective mixing is much smaller than for non-rotating flows. In our simulations, we calculate Ro using the r.m.s. velocity of the fluid in each convective layer as a proxy for $U_{\mathrm{conv}}$, and set $\ell$ to the thickness of a convective layer in the staircase. Figure 4 shows profiles of the r.m.s. velocity for the rotating case, at different stages during the evolution of the staircase. As expected, we find that as layers merge, the freefall distance of each surviving layer increases, leading to an increase in the convective velocity (see Equation 2 for how $U_{\mathrm{conv}}$ depends on $\ell$). However, the ratio $U_{\mathrm{conv}}(\ell)/\ell$ decreases. Consequently, the Rossby number, shown at the top of each curve in Figure 4, decreases after each merger, starting with $\mathrm{Ro}\approx 0.107$ in the 6-layers phase, and ultimately reaching $\mathrm{Ro}\approx 0.044$. We also note that initially, when there are 6 layers in the fluid, the velocity and size of each layer is uniform. Later on, the outer convective layers have a larger speed than the interior layers, engulfing them and leaving only two layers that slowly erode each other until a single convective layer remains.
@@ 3 Analytic model for the erosion phase
A complete characterization of all the timescales involved in the problem, and how they change with the choice of parameters is beyond the scope of this work. However, in what follows, we present a detailed model for the erosion stage that reproduces our numerical results.
Figure 5: Panel (a): Density profile at the beginning of the erosion phase for the non-rotating and rotating cases. Note that in the rotating case, the density is not constant within the convection zone. Instead, there is a positive density gradient due to the thermal structure of the fluid in the convection zone (in nonrotating convection, the temperature becomes isothermal whereas in rotating convection, the fluid develops a small temperature gradient). Panel (b): Evolution of the density contrast between the two layers during the erosion phase. For both cases $\Delta\rho$ decreases linearly with time, but the rotating case decays at a slower rate. Further, the initial density contrast at the beginning of the erosion phase is smaller for the rotating case (likely due to diffusion since it takes longer for the fluid to reach this stage).
@@
where the preceding factor of 2 in the final expression accounts for entrainment from above and below the interface, $\rho_{0}$ is the average density of the fluid, and $\varepsilon$ is the mixing efficiency (0.5absent0.5$\approx$ 0.5≈ 0.51111, e.g., Fuentes & Cumming, 2020).
We see that given $U_{\mathrm{conv}}$, Equation (1) can be integrated to obtain $\Delta\rho(t)$. In the non-rotating and rapidly-rotating limits, typical convective speeds ($U_{\rm NR}$ and $U_{\rm R}$, respectively) can be estimated using mixing-length theory and turbulent, diffusion-free scaling laws for convective heat transfer (e.g., Stevenson, 1979; Ingersoll & Pollard, 1982; Barker et al., 2014; Aurnou et al., 2020). Following the derivation of Fuentes et al. (2023), we obtain
@@
where the numbers were inferred from Cumming et al. (2018) and Helled et al. (2022). This is a difference of two orders of magnitude, and the mixing time is longer than the current age of Jupiter or Saturn. We emphasize that evaluating the numbers at later times would give even longer timescales, since convection becomes weaker as the planet’s luminosity decreases with time.
@@ 4 Discussion
Despite the simplifications made in our model (e.g., ignoring spherical geometry and density stratification) and the uncertainties in the interior parameters of gas giants (particularly for $\Delta\rho_{0}$ and the mixing efficiency $\varepsilon$), we have demonstrated that the effect of rotation is substantial in rapidly rotating gas giants, increasing the erosion time of the staircase by two orders of magnitude, for Jovian conditions at $t\sim 10^{8}~{}\mathrm{yrs}$ after formation. The effect is even larger at later times since the convective strength becomes smaller as the planet’s luminosity decreases with time. Also, a smaller value of the mixing efficiency for the rotating case will just make the timescale longer, once again supporting our hypothesis that rotation prolongs the lifetime of a convective staircase.
Characterizing the timescale of the initial mergers is also of great interest, as it would allow us to determine the current evolutionary stage of the fuzzy cores in Jupiter and Saturn. For example, Saturn’s seismology indicates the presence of a stably stratified region, which likely requires a semi-convective staircase with many steps, giving the planet a dense spectrum of gravity modes. If Saturn were in the erosion phase, there would be only a single interface and hence a single interface mode in the pulsation spectrum, which contradicts the ring-seismology inferences (Fuller, 2014; Belyaev et al., 2015; Mankovich & Fuller, 2021). Unlike the erosion phase, where the staircase consists of only two steps of equal size throughout the process, the initial mergers occur in a fluid region with many steps of size $h$, which merge over time, causing $h$ to increase. From Equations (6) and (7) in Fuentes et al. (2023), and using the same parameters, the growth rate of a convective layer by engulfing fluid from a stable region below is $h_{\mathrm{NR}}\sim 3\times 10^{9}~{}\mathrm{cm}~{}(t/4.5~{}\mathrm{Gyr})^{1/2}$ for the non-rotating case, and $h_{\mathrm{R}}\sim 2\times 10^{8}~{}\mathrm{cm}~{}(t/4.5~{}\mathrm{Gyr})^{5/12}$ for the rotating case. This suggests that over the age of the solar system, the planet would be nearly fully mixed in the non-rotating case, while about $\sim 10$ convective layers would remain in the staircase for the rotating case.
Another factor that could influence convective mixing and the lifespan of a staircase is the presence of magnetic fields. For example, the preference for fluid flows to move parallel to the magnetic field direction means that a vertical field geometry can significantly reduce horizontal mixing when the field strength is sufficiently large (e.g., Yan et al., 2019). Recently, Sanghi et al. (2022) investigated semiconvection in the presence of a uniform vertical background magnetic field, finding that a sufficiently strong field delays layer formation and substantially reduces the internal flux across the layers. However, different field geometries could lead to different results (e.g., Calkins et al., 2023).
@@
We encourage future work to either confirm or refute our findings using a larger suite of numerical simulations. More detailed models of the growth and merging rates of semiconvective layers would significantly improve the estimates mentioned above and help assess the possible structures of giant planets over evolutionary timescales. It is also important to emphasize that semiconvection is not only relevant for the thermal evolution and transport processes in giant planets, but also in stars (e.g., Spruit, 2013; Zaussinger & Spruit, 2013; Ding & Li, 2014; Moore & Garaud, 2016). Recently, semiconvection has been shown to be crucial for understanding the internal dynamics, magnetism, and potential habitability of icy moons (e.g., Vance et al., 2021; Wong et al., 2022; Naseem et al., 2023; Idini & Nimmo, 2024). The Rossby number in icy moons is also small ($\mathrm{Ro}\ll 1$, see, e.g., Bire et al., 2022) and therefore the results obtained in this work could be applicable to them.
@@
We thank Keith Julien for the countless conversations on fluid dynamics, in particular the ones on rotationally-constrained turbulence. You were not only a great scientist, but also a good friend, with an infectious smile. We also thank Pascale Garaud and Jim Fuller for useful conversations on double-diffusive convection and about the application of our results to Saturn and Jupiter. J.R.F. is supported by NASA through the Solar System Workings grant 80NSSC24K0927, and Heliophysics grants 80NSSC19K0267 and 80NSSC20K0193. B.W.H. is supported by NASA Heliophysics through grants 80NSSC20K0193, 80NSSC24K0125, and 80NSSC24K0271. A.F. acknowledges support from NASA HTMS grant 80NSSC20K1280, and from the George Ellery Hale Postdoctoral Fellowship in Solar, Stellar and Space Physics at CU Boulder. E.H.A. acknowledges support from NSF grant PHY-2309135 and Simons Foundation grant (216179, LB) to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Fluid equations and numerical methods
We express the fluid quantities as the sum of a linear hydrostatic background (denoted by the subscript 0) and a dynamic perturbation to the background (denoted by primes), e.g., the total temperature and composition are expressed as $T=T_{0}(z)+T^{\prime}$, and $C=C_{0}(z)+C^{\prime}$, respectively. The density perturbations satisfy $\rho^{\prime}/\rho_{0}\ll 1$, and are related to $T^{\prime}$ and $C^{\prime}$ through $\rho^{\prime}=\rho_{0}(\beta C^{\prime}-\alpha T^{\prime})$, as demanded by the Boussinesq approximation (Spiegel & Veronis, 1960). Here, $\beta$ and $\alpha$ are the coefficients of compositional and thermal contraction/expansion (both assumed positive constants), respectively. Before presenting the fluid equations, we non-dimensionalize them using $[T]=|\partial_{z}T_{0}|L_{z}$, $[C]=|\partial_{z}C_{0}|L_{z}$ as units of temperature and composition. We use the domain’s depth $L_{z}$ as the unit of length, and the thermal diffusion time $\tau_{T}=L^{2}_{z}/\kappa_{T}$ (where $\kappa_{T}$ is the thermal diffusivity) as the unit of time. By this choice, a unit of pressure corresponds to $[P]=\rho_{0}(\kappa_{T}/L_{z})^{2}$. The dimensionless equations are